OpenStudy (anonymous):

find a, b , and h so that f(x)= a sin (b(x-h)) f(x)=5cosx+2sinx

OpenStudy (anonymous):

how do you solve this step by step :)

OpenStudy (anonymous):

what about b and h those are confusing :/

OpenStudy (anonymous):

sorry that was wrong \[a=\sqrt{5^2+2^2}\]

OpenStudy (anonymous):

forget the "b"

OpenStudy (anonymous):

it is \[a\cos(x)+b\sin(x)=\sqrt{a^2+b^2}\sin\left(x+\tan^{-1}(\frac{a}{b})\right)\]

OpenStudy (anonymous):

thats how you find h?

OpenStudy (anonymous):

yes, that and a calculator in this case

OpenStudy (anonymous):

umm is that the formula every time? :)

OpenStudy (anonymous):

i put in the calc but what #'s am i plugging in?